Please do each step. Show your work.
Solve each aequation for x.
Part 1: log[3](x+5)=2
Step 1: Write the given logarithmic equation in exponential form.
Step 2: Solve the exponential equation from step 1 for x.
Step 3: Check for extraneous solutions.
Step 4: State the solution for the given equation.

Answer by nerdybill(7384)

Solve each aequation for x.
Part 1: log[3](x+5)=2
Step 1: Write the given logarithmic equation in exponential form.
x+5 = 3^2
Step 2: Solve the exponential equation from step 1 for x.
x+5 = 3^2
x+5 = 9
x = 4
Step 3: Check for extraneous solutions.
None.
Step 4: State the solution for the given equation.
x = 4